OpenStudy (anonymous):

integral of (x/(x+1)

ganeshie8 (ganeshie8):

hint : x = x + 1 - 1

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ x }{ x+1 }dx\]

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ x + 1 - 1 }{ x+1 }\]

OpenStudy (anonymous):

thank you both

OpenStudy (anonymous):

No problem!